NEED HELP ASAP. CAN SOMEONE HELP ME WITH THIS WORD PROBLEM?
SUPPOSE THAT THE FUNCTION P =12 + 50ln x respresents the percentage of inbound e-mail in the U.S that is considered spam. where x is the number of years after 2001. Carry all calculation to six decimals on each step, when necessary.
a)USE THIS MODEL TO APPROXIMATE THE PERCENTAGE OF SPAM IN THE YEAR 2005.
ANSWER:
SHOW YOUR WORK:
B)USE THIS MODEL TO APPROXIMATE THE YEAR THE PERCENT OF SPAM WILL REACH 90% PROVIDED THAT LAW ENFORCMENT REGARDING SPAMMERS DOES NOT CHANGE?
ANSWER:
SHOW YOUR WORK PLEASE:
THANK YOU VERY MUCH TUTOR

Answer by stanbon(75887)

P =12 + 50ln x respresents the percentage of inbound e-mail in the U.S that is considered spam. where x is the number of years after 2001. Carry all calculation to six decimals on each step, when necessary.
-------------------
a)USE THIS MODEL TO APPROXIMATE THE PERCENTAGE OF SPAM IN THE YEAR 2005.
ANSWER:
SHOW YOUR WORK:
x = 2005-2001 = 4
P = 12 + 50(ln(4))
P = 12 + 69.314718 = 71.314718%
=========================
B)USE THIS MODEL TO APPROXIMATE THE YEAR THE PERCENT OF SPAM WILL REACH 90% PROVIDED THAT LAW ENFORCMENT REGARDING SPAMMERS DOES NOT CHANGE?
ANSWER:
90 = 12 + 50ln(x)
78 = 50 ln(x)
ln(x) = 38/25
x = e^(38/25)
x = 4.572225 yrs
----------------
2001 + 4.572225 = 2005.572225 or in the year 2006
====================
Cheers,
Stan H.